OpenStudy (anonymous):

1. x²+4=0 2. x²-6x+1=0 Solve each equation, write irrational numbers in simplest radical form. Describe the strategy you used to get your solution and tell why you chose that strategy.

OpenStudy (anonymous):

1) subtract 4 from both sides. Remember i=sqrt(-1)

OpenStudy (anonymous):

2) use the quadratic formula

OpenStudy (anonymous):

go to youtube and search for "Quadratic formula song". Learn it. Memorize it. Appriciate it!

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

so, would it be x²=-4?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

x²-6x+1=0 +6 +6 x²+1=6 -1 -1 x²=5 is this correct?

OpenStudy (anonymous):

you dropped an x there x²+1=6x

OpenStudy (anonymous):

just use the quadratic formula, it's much easier